OpenStudy (anonymous):

Integrate the following using partial fractions indefinite integral of (22−(6t2+11t))/((t+4)(t−1)(t−2)) the partial fraction form is -(1/(t+4))+-(1/(t-1))-(4/(t-2)) the only problem i have is that i keep getting - ln(t-1)-4( ln(t-2))-ln(t+4)+C, but the program im using says it's wrong

zepdrix (zepdrix):

\[\large \int\limits \frac{-1}{t+4}+\frac{-1}{t-1}+\frac{-4}{t-2}\quad dt\] \[\large -\ln(t+4)-\ln(t-1)-4\ln(t-2)+C\]Program says it's wrong? :( Hmmm it looks correct...

OpenStudy (anonymous):

I'll email my instructors then, thanks anyway :)